# Probabilistic modeling of Cherenkov emission from particle showers

### v1

@@ 3.2 Parametric model of the amplitude
In refs. Raedel2012; Radel:2012ij distributions of $\hat{\ell}_{\text{tot}}$ from Geant4 were fit to a normal distribution at each energy. Two examples are shown in the right panel figure˜1 for $1\text{\,}\mathrm{TeV}$ $e^{-}$ and $\pi^{+}$ as dashed and dotted black lines, respectively. For comparison, the corresponding FLUKA $e^{-}$ ($\pi^{+}$) MC distribution is shown in blue (orange) and exhibits left (right) skew that deviates from the normal distribution. Additional nuclear effects which were not enabled in the prior Geant4 simulations but were in FLUKA lead to a noticeable tail in the $\hat{\ell}_{\text{tot}}$ distribution of EM showers. The right panel of figure˜2 shows their energy dependence, highlighting that the skew is consistent across different energies.
In order to more accurately model FLUKA $\hat{\ell}_{\text{tot}}$ distributions, and due to the differing skewness, new functional forms for EM and hadronic showers were used. The $\hat{\ell}_{\text{tot}}$ distribution from hadron primaries are modeled by a skew normal distribution Azzalini1999Statistical, while $e^{-}$ and $\gamma$ primaries rely on the normal-inverse Gaussian (NIG) distribution 10.1111/1467-9469.00045. Both are implemented in the SciPy toolkit 2020SciPy-NMeth, from which their definitions are adapted. The skew normal is a three parameter distribution with shape, $\alpha$, location, $\xi$, and scale, $\omega$, defined in terms of $z\equiv(x-\xi)/\omega$ as
$\mathrm{SN}(x;\alpha,\xi,\omega)=\frac{2\phi(z)\Phi(\alpha z)}{\omega}$ (5)
@@
Figure 6: The FLUKA $\hat{\ell}_{\text{tot}}$ distributions in the right panel are identical to those in that of figure˜1. A fit to a normal-inverse Gaussian (skew normal) distribution for $e^{-}$ ($\pi^{+}$) results in the probability density function indicated by the dashed blue (dotted orange) line. Both are improved descriptions of the data compared to the black dashed and dotted lines in figure˜1 from refs. Raedel2012; Radel:2012ij. The left panel shows fitted parameters as a function of the primary $e^{-}$ ($\pi^{+}$) energy in blue (orange) using different markers. For $e^{-}$ the parameters enter eq.˜6, and $\pi^{+}$ eq.˜5. Predictions from $g_{p}(E;\bm{\hat{t}}_{p},s_{p})$ are shown as corresponding lines. See text for more details.
@@
To illustrate, the right panel of figure˜6 shows the NIG (SN) PDF in dashed blue (dotted orange) obtained from eq.˜7 for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$). The underlying $\hat{\ell}_{\text{tot}}$ distributions from FLUKA are shown as well, and are identical to that of the right panel of figure˜1. Agreement between the PDFs and underlying distributions are noticeably improved. The left panel shows fitted parameter values at discrete energies using eq.˜6 for $e^{-}$ (blue) and eq.˜5 for $\pi^{+}$ (orange). Shape parameters are indicated by triangles, location by circles, and scale by squares. Lines that appear to interpolate the markers correspond to predictions from eq.˜7. To reasonable approximation $\xi$ scales linearly with energy Raedel2012; Radel:2012ij. Deviations from linearity can be seen by comparing $\xi$ between the blue and orange, and in order to accurately model all parameters a sixth degree polynomial in $\log_{10}E$ is fitted according to eq.˜8. This procedure is repeated for all particles listed in section˜2.1.
@@ 3.3 Accuracy
The gamma distribution was successfully fitted to almost every simulated shower. At most a single fit failed out of $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}\mathrm{0}$ simulations at a given energy, mostly below $1\text{\,}\mathrm{TeV}$, and no more than four over all energies for any given primary. To quantify the accuracy, every gamma distribution was compared against its FLUKA simulation and the Wasserstein distance, $W_{1}=\int_{-\infty}^{\infty}|\tilde{F}(x)-F(x)|dx$, was computed 10.1561/2200000073. Here, $\tilde{F}(x)$ is the sum of the binned FLUKA $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ up to distance $x$, and $F(x)$ the CDF of its fitted gamma distribution. Note that each simulation thus corresponds to an independent data point. For each simulated energy, the resulting $W_{1}$ $50\text{\,}\%$ and $95\text{\,}\%$ quantiles are shown as colored lines in the top three panels of figure˜8 for all primaries, listed in the top-left legend. For comparison, the median obtained assuming the average-shape parametrizations from refs. Raedel2012; Radel:2012ij are included as gray lines. A significantly improved characterization of the shape is observed with the individually fitted gamma distributions. The metric $W_{1}$ was chosen, as opposed to a statistic, since the intent is not to claim that simulated shower shapes correspond to the gamma distribution; as discussed earlier, the gamma distribution is a simple approximation. Instead, $W_{1}$ is a measure of proper distance between two distributions, and highlights the energy dependence and improved agreement relative to refs. Raedel2012; Radel:2012ij.
Figure 8: The colored lines in the top three panels shows $50\text{\,}\%$ and $95\text{\,}\%$ quantiles of the Wasserstein distance, $W_{1}$, as a function of energy for all primaries. Additionally, the median obtained assuming the average parametrizations from refs. Raedel2012; Radel:2012ij are included as gray lines. From left to right, the panels correspond to EM, mesonic, and baryonic primaries, as indicated by the top-left legend. The bottom three panels show the KS statistic for $a^{\prime}$, $b^{\prime}$ and $\hat{\ell}_{\text{tot}}$ also as a function of energy for all primaries. The different primaries are shown as different markers, listed in the top-right legend. For each primary, at a given energy, the KS statistic was computed by comparing $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}$ samples from the model against the values calculated directly from the $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}\mathrm{0}$ FLUKA simulations.
@@ 4 Application to high-energy neutrino interactions
The models described in section˜3 are for a single primary particle. In collisions where multiple final state particles are produced, the model can be applied individually to each one. Here, we use PYTHIA8 Sjostrand:2014zea to simulate electron-neutrino DIS interactions off a proton target. Final state particles with a proper lifetime of $c\tau_{0}>$0.5\text{\,}\mathrm{mm}$\$ are stored directly. Particles with $c\tau_{0}<$0.5\text{\,}\mathrm{mm}$\$, which includes most short-lived particles like $D$-mesons and $B$-mesons, are propagated and decayed by PYTHIA. Longer-lived decay products are then stored. For heavy mesons, at energies above a few tens of TeV the interplay of decay and interaction lengths makes their parametrization challenging, and a MC approach is necessary for accurate simulation. Further assuming that antiparticles can use the same parametrizations as their particle counterparts, the final list of particles and their energies can be used to individually sample $\text{d}\hat{\ell}/\text{d}x$.
@@ 5 Discussion
Using sophisticated particle simulation software like FLUKA Ferrari:2005zk; Ballarini:2024isa, the most subtle details of particle interactions can be modeled to high accuracy. In applications to high-energy Cherenkov calorimeters, such as neutrino telescopes, it is essential to model detectable signatures that are relevant for analyses. While existing models have approximated the energy deposition characteristics in particle showers, the method outlined in this paper more accurately captures the shape and scale of the Cherenkov yield. A deterministic model cannot represent variations in the position and dispersion of a shower Radel:2012ij; Raedel2012. With the addition of relatively minor computational complexity we can capture both of these factors, as detailed in section˜3.1 and section˜3.2.
@@
A simple-to-use Python package containing all the models is available acem. The best-fit B-spline coefficients, $\bm{\hat{\theta}}$, their defining knots, as well as the polynomial coefficients, $\bm{\hat{t}}_{p}$, for modeling $\hat{\ell}_{\text{tot}}$ are saved in the NumPy format harris2020array. As these are simply arrays of floats, they can be ported to other frameworks and languages in a relatively straightforward manner using existing B-spline libraries Whitehorn:2013nh.
@@
Existing simulation of high-energy neutrino interactions at neutrino telescopes such as IceCube does not incorporate shower-to-shower fluctuations, and instead rely on the average profiles from refs. Radel:2012ij; Raedel2012. When such fluctuations are included, they can lead to shower profiles that differ substantially from the mean, and can modify the Cherenkov light yield of neutrino interactions as shown in section˜4. Individual showers that peak at upwards of $10\text{\,}\mathrm{m}$ from the interaction vertex are possible. For analyses searching for rare signals based on small separation distances, it becomes more important to accurately simulate in-medium particle showers. The models presented in section˜3 are a step towards more accurate representations of the Cherenkov yield from high-energy particle showers in ice and water, and may be of use in simulations for existing and upcoming experiments.
@@ Appendix A Additional enabled physics in FLUKA
Additional physics settings enabled in the FLUKA simulations described in section˜2.1 are given below. For a description of their effects see ref. Ballarini:2024isa and the FLUKA online manual at https://www.fluka.eu.
@@ Appendix B Smoothing parameters
Here, dev is the deviance, a metric of the closeness between the data and the model, and dim is the effective dimension of the model space, or the degrees of freedom we had available in fitting. This is calculated as the trace of the hat-matrix for the general linear model 10.1214/ss/1038425655.
@@ Appendix D Additional results for $100\text{\,}\mathrm{TeV}$ electron neutrino DIS
Figure˜10 shows $\mathrm{8}\mathrm{0}$ $\text{d}\hat{\ell}/\text{d}x$ profiles of $100\text{\,}\mathrm{TeV}$ electron-neutrino interactions, as discussed in section˜4. The first two correspond to that shown in figure˜9. Each neutrino interacts off a fixed proton target, with hadronization and decays of particles up to $c\tau_{0}<$0.5\text{\,}\mathrm{mm}$\$ simulated with PYTHIA8 Sjostrand:2014zea. The resulting final state $e^{-}$, $\gamma$ and hadrons are passed to the model of section˜3, which is used to sample $\text{d}\hat{\ell}/\text{d}x$ according to their energies. In figure˜10, panels highlighted in blue correspond to NC interactions, otherwise CC.
Figure 10: Eighty $\text{d}\hat{\ell}/\text{d}x$ results based on $100\text{\,}\mathrm{TeV}$ electron-neutrino interactions generated with PYTHIA8 Sjostrand:2014zea. The first two correspond to that of figure˜9. Panels highlighted in blue correspond to NC interactions, otherwise CC. Within each panel, blue (orange) dashed lines show sampled $\text{d}\hat{\ell}/\text{d}x$ distributions for EM (hadron) primaries, and solid lines show the sum for each category. The total is shown in black. To give a sense of both the scale and shape a logarithmic scale is used in the y-axis.

### v2

@@ 3.2 Parametric model of the amplitude
In refs. Rädel (2012); Radel and Wiebusch (2013) distributions of $\hat{\ell}_{\text{tot}}$ from Geant4 were fit to a normal distribution at each energy. Two examples are shown in the right panel figure˜1 for $1\text{\,}\mathrm{TeV}$ $e^{-}$ and $\pi^{+}$ as dashed and dotted black lines, respectively. For comparison, the corresponding FLUKA $e^{-}$ ($\pi^{+}$) MC distribution is shown in blue (orange) and exhibits left (right) skew that deviates from the normal distribution. Additional nuclear effects which were not enabled in the prior Geant4 simulations but were in FLUKA lead to a noticeable tail in the $\hat{\ell}_{\text{tot}}$ distribution of EM showers. The right panel of figure˜2 shows their energy dependence, highlighting that the skew is consistent across different energies.
In order to more accurately model FLUKA $\hat{\ell}_{\text{tot}}$ distributions, and due to the differing skewness, new functional forms for EM and hadronic showers were used. The $\hat{\ell}_{\text{tot}}$ distribution from hadron primaries are modeled by a skew normal distribution Azzalini and Capitanio (1999), while $e^{-}$ and $\gamma$ primaries rely on the normal-inverse Gaussian (NIG) distribution Barndorff-Nielsen (1997). Both are implemented in the SciPy toolkit Virtanen et al. (2020), from which their definitions are adapted. The skew normal is a three parameter distribution with shape, $\alpha$, location, $\xi$, and scale, $\omega$, defined in terms of $z\equiv(x-\xi)/\omega$ as
$\mathrm{SN}(x;\alpha,\xi,\omega)=\frac{2\phi(z)\Phi(\alpha z)}{\omega},$ (5)
@@
Figure 6: The FLUKA $\hat{\ell}_{\text{tot}}$ distributions in the right panel are identical to those in that of figure˜1. A fit to a normal-inverse Gaussian (skew normal) distribution for $e^{-}$ ($\pi^{+}$) results in the probability density function indicated by the dashed blue (dotted orange) line. Both are improved descriptions of the data compared to the black dashed and dotted lines in figure˜1 from refs. Rädel (2012); Radel and Wiebusch (2013). The left panel shows fitted parameters as a function of the primary $e^{-}$ ($\pi^{+}$) energy in blue (orange) using different markers. For $e^{-}$ the parameters enter eq.˜6, and $\pi^{+}$ eq.˜5. Predictions from $g_{p}(E;\bm{\hat{t}}_{p},s_{p})$ are shown as corresponding lines. See text for more details.
@@
To illustrate, the right panel of figure˜6 shows the NIG (SN) PDF in dashed blue (dotted orange) obtained from eq.˜7 for a $1\text{\,}\mathrm{TeV}$ $e^{-}$ ($\pi^{+}$). The underlying $\hat{\ell}_{\text{tot}}$ distributions from FLUKA are shown as well, and are identical to that of the right panel of figure˜1. Agreement between the PDFs and underlying distributions are noticeably improved. The left panel shows fitted parameter values at discrete energies using eq.˜6 for $e^{-}$ (blue) and eq.˜5 for $\pi^{+}$ (orange). Shape parameters are indicated by triangles, location by circles, and scale by squares. Lines that appear to interpolate the markers correspond to predictions from eq.˜7. To reasonable approximation $\xi$ scales linearly with energy Rädel (2012); Radel and Wiebusch (2013). Deviations from linearity can be seen by comparing $\xi$ between the blue and orange, and in order to accurately model all parameters a sixth degree polynomial in $\log_{10}E$ is fitted according to eq.˜8. This procedure is repeated for all particles listed in section˜2.1.
@@ 3.3 Accuracy
The gamma distribution was successfully fitted to almost every simulated shower. At most a single fit failed out of $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}\mathrm{0}$ simulations at a given energy, mostly below $1\text{\,}\mathrm{TeV}$, and no more than four over all energies for any given primary. To quantify the accuracy, every gamma distribution was compared against its FLUKA simulation and the Wasserstein distance, $W_{1}=\int_{-\infty}^{\infty}|\tilde{F}(x)-F(x)|dx$, was computed Peyré and Cuturi (2019). Here, $\tilde{F}(x)$ is the sum of the binned FLUKA $\hat{\ell}_{\text{tot}}^{-1}\text{d}\hat{\ell}/\text{d}x$ up to distance $x$, and $F(x)$ the CDF of its fitted gamma distribution. Note that each simulation thus corresponds to an independent data point. For each simulated energy, the resulting $W_{1}$ $50\text{\,}\%$ and $95\text{\,}\%$ quantiles are shown as colored lines in the top three panels of figure˜8 for all primaries, listed in the top-left legend. For comparison, the median obtained assuming the average-shape parametrizations from refs. Rädel (2012); Radel and Wiebusch (2013) are included as gray lines. A significantly improved characterization of the shape is observed with the individually fitted gamma distributions. The metric $W_{1}$ was chosen, as opposed to a statistic, since the intent is not to claim that simulated shower shapes correspond to the gamma distribution; as discussed earlier, the gamma distribution is a simple approximation. Instead, $W_{1}$ is a measure of proper distance between two distributions, and highlights the energy dependence and improved agreement relative to refs. Rädel (2012); Radel and Wiebusch (2013).
Figure 8: The colored lines in the top three panels shows $50\text{\,}\%$ and $95\text{\,}\%$ quantiles of the Wasserstein distance, $W_{1}$, as a function of energy for all primaries. Additionally, the median obtained assuming the average parametrizations from refs. Rädel (2012); Radel and Wiebusch (2013) are included as gray lines. From left to right, the panels correspond to EM, mesonic, and baryonic primaries, as indicated by the top-left legend. The bottom three panels show the KS statistic for $a^{\prime}$, $b^{\prime}$ and $\hat{\ell}_{\text{tot}}$ also as a function of energy for all primaries. The different primaries are shown as different markers, listed in the top-right legend. For each primary, at a given energy, the KS statistic was computed by comparing $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}$ samples from the model against the values calculated directly from the $\mathrm{1}\mathrm{0}\mathrm{0}\mathrm{0}\mathrm{0}$ FLUKA simulations.
@@ 4 Application to high-energy neutrino interactions
The models described in section˜3 are for a single primary particle. In collisions where multiple final state particles are produced, the model can be applied individually to each one. Here, we use PYTHIA8 Sjöstrand et al. (2015) to simulate electron-neutrino DIS interactions off a proton target. Final state particles with a proper lifetime of $c\tau_{0}>$0.5\text{\,}\mathrm{mm}$\$ are stored directly. Particles with $c\tau_{0}<$0.5\text{\,}\mathrm{mm}$\$, which includes most short-lived particles like $D$-mesons and $B$-mesons, are propagated and decayed by PYTHIA. Longer-lived decay products are then stored. For heavy mesons, at energies above a few tens of TeV the interplay of decay and interaction lengths makes their parametrization challenging, and a MC approach is necessary for accurate simulation. Further assuming that antiparticles can use the same parametrizations as their particle counterparts, the final list of particles and their energies can be used to individually sample $\text{d}\hat{\ell}/\text{d}x$.
@@ 5 Discussion
Using sophisticated particle simulation software like FLUKA Ferrari et al. (2005); Ballarini and others (2024), the most subtle details of particle interactions can be modeled to high accuracy. In applications to high-energy Cherenkov calorimeters, such as neutrino telescopes, it is essential to model detectable signatures that are relevant for analyses. While existing models have approximated the energy deposition characteristics in particle showers, the method outlined in this paper more accurately captures the shape and scale of the Cherenkov yield. A deterministic model cannot represent variations in the position and dispersion of a shower Radel and Wiebusch (2013); Rädel (2012). With the addition of relatively minor computational complexity we can capture both of these factors, as detailed in section˜3.1 and section˜3.2.
@@
A simple-to-use Python package containing all the models is available Crawshaw and Yuan (2025). The best-fit B-spline coefficients, $\bm{\hat{\theta}}$, their defining knots, as well as the polynomial coefficients, $\bm{\hat{t}}_{p}$, for modeling $\hat{\ell}_{\text{tot}}$ are saved in the NumPy format Harris et al. (2020). As these are simply arrays of floats, they can be ported to other frameworks and languages in a relatively straightforward manner using existing B-spline libraries Whitehorn et al. (2013).
@@
Existing simulation of high-energy neutrino interactions at neutrino telescopes such as IceCube does not incorporate shower-to-shower fluctuations, and instead rely on the average profiles from refs. Radel and Wiebusch (2013); Rädel (2012). When such fluctuations are included, they can lead to shower profiles that differ substantially from the mean, and can modify the Cherenkov light yield of neutrino interactions as shown in section˜4. Individual showers that peak at upwards of $10\text{\,}\mathrm{m}$ from the interaction vertex are possible. For analyses searching for rare signals based on small separation distances, it becomes more important to accurately simulate in-medium particle showers. The models presented in section˜3 are a step towards more accurate representations of the Cherenkov yield from high-energy particle showers in ice and water, and may be of use in simulations for existing and upcoming experiments.
@@ Appendix A Additional enabled physics in FLUKA
Additional physics settings enabled in the FLUKA simulations described in section˜2.1 are given below. For a description of their effects see ref. Ballarini and others (2024) and the FLUKA online manual at https://www.fluka.eu.
@@ Appendix B Smoothing parameters
Here, dev is the deviance, a metric of the closeness between the data and the model, and dim is the effective dimension of the model space, or the degrees of freedom we had available in fitting. This is calculated as the trace of the hat-matrix for the general linear model Eilers and Marx (1996).
@@ Appendix D Additional results for $100\text{\,}\mathrm{TeV}$ electron neutrino DIS
Figure˜10 shows $\mathrm{8}\mathrm{0}$ $\text{d}\hat{\ell}/\text{d}x$ profiles of $100\text{\,}\mathrm{TeV}$ electron-neutrino interactions, as discussed in section˜4. The first two correspond to that shown in figure˜9. Each neutrino interacts off a fixed proton target, with hadronization and decays of particles up to $c\tau_{0}<$0.5\text{\,}\mathrm{mm}$\$ simulated with PYTHIA8 Sjöstrand et al. (2015). The resulting final state $e^{-}$, $\gamma$ and hadrons are passed to the model of section˜3, which is used to sample $\text{d}\hat{\ell}/\text{d}x$ according to their energies. In figure˜10, panels highlighted in blue correspond to NC interactions, otherwise CC.
Figure 10: Eighty $\text{d}\hat{\ell}/\text{d}x$ results based on $100\text{\,}\mathrm{TeV}$ electron-neutrino interactions generated with PYTHIA8 Sjöstrand et al. (2015). The first two correspond to that of figure˜9. Panels highlighted in blue correspond to NC interactions, otherwise CC. Within each panel, blue (orange) dashed lines show sampled $\text{d}\hat{\ell}/\text{d}x$ distributions for EM (hadron) primaries, and solid lines show the sum for each category. The total is shown in black. To give a sense of both the scale and shape a logarithmic scale is used in the y-axis.